# A Portable Parton-Level Event Generator for the High-Luminosity LHC FERMILAB-PUB-23-641-T, MCNET-23-18 A Portable Parton-Level Event Generator

### v3

@@ 5 Performance
We note that the core algorithms of Pepper (then referred to as Blockgen) have been studied in detail in our gluon-only pathfinder study [32], and we refer the reader to this study for more technical performance results and comparisons with alternative algorithms. In the following, we instead focus on portability and performance comparisons on different architectures. We only repeat here that the performance of Pepper is still memory bound on the GPU. In fact, the addition of fermions in Pepper requires the use of complex-valued currents in the recursion, which further intensifies memory operations.
@@ 5.1 Baseline CPU performance
Figure 3 presents a comparison of the CPU version of Pepper and the parton-level event generator Comix 666We use a non-default mode for Comix, splitting the process into different Monte-Carlo channels, grouped by gluons, valence and sea quarks. This mode of operation is similar to Pepper and increases the efficiency of the event generation. [81]. We measure the rate at which unweighted events are generated and stored for and production at the LHC. The center-of-mass energy of the collision is set to TeV, and the partons are required to fulfill GeV, and . For production, an additional cut of is applied to the invariant mass of the dilepton system. For simplicity, the renormalization and factorization scales are fixed to in production, and to in production. For better comparability of the measurements, we show Comix both in combination with its recursive phase space generator [81] and with the modified basic Chili integrator [20] that is also used by Pepper, cf. Sec. 2 for details. The raw data for Fig. 3 are listed in App. A.
@@ 5.2 Performance on different hardware
Using a portability framework like Kokkos in the Pepper event generator is a novel feature for a production-ready parton-level event generator, and for much of the HEP software stack in general. We have therefore tested the performance of the Kokkos Pepper variant against both a native CUDA and a native single-threaded CPU implementation, and looked at results on an A100 GPU and an Intel Core i3-8300 CPU at 3.70 GHz CPU. When comparing the event throughput of the Kokkos and native CUDA implementations on the A100 GPU, we find that the performance of the Kokkos variant agrees within a factor of two, with the performance gap disappearing as computational complexity increases, i.e. with increasing process multiplicity. Generally, the native algorithm outperforms the Kokkos version. This result establishes that using a portability framework is possible without compromising significantly on performance when using a GPU, while at the same time giving access to a much wider range of architectures and computing paradigms. With that, we only show Kokkos result in the following. We note, however, that our current Kokkos implementation on the CPU does not utilize vectorization capabilities, while our native C++ implementation does so with the help of the VCL library [90], which provides significant speedups on the CPU. Because the focus of the analysis here is on portability, these improvements are not yet included in Figs. 4 and 5. We expect to provide a Kokkos implementation with vectorization capabilities in the future.
@@ Appendix C Parton-level validation
The comparison results for the two Z+jets observables are shown in Fig. 8. For details on the presentation and the observables, we refer to Sec. 4. The agreement is again quantified using a Kolmogorow–Smirnov test. The -values for each jet multiplicity can be found on the plot. They are all greater than our chosen confidence level of , i.e. we do not reject the null hypothesis of the two underlying distributions being identical.
Finally, the comparison results for the two +jets observables are shown in Fig. 9. Again, the results of a Kolmogorow–Smirnov test are quoted on the plot, with all -values being greater than 0.05. Thus, also in this case, we do not reject the null hypothesis.
@@ Appendix D Data layouts and CPU performance
The two approaches yield different caching opportunities. In Fig. 10, we compare their performance on an Apple M2 Pro chip for different event batch sizes for the jets process. The performance is measured in events per hour, normalized to the SoA event rate for a batch size of one. We find small performance improvements of about when increasing the batch size to 10…100 events, with the SoA performance being within a few percent with the AoS performance, indicating that good cache efficiency is achieved for both data layouts.

### v4

@@ 5 Performance
We note that the core algorithms of Pepper (then referred to as Blockgen) have been studied in detail in our gluon-only pathfinder study [32], and we refer the reader to this study for more technical performance results and comparisons with alternative algorithms. In addition, we provide Pepper profiling data for () and () production runs on an Nvidia V100 GPU for further study [81, 82]. However, in the following, we instead focus on high-level portability and performance comparisons on different architectures. We only repeat here that the performance of Pepper is still memory bound on the GPU. In fact, the addition of fermions in Pepper requires the use of complex-valued currents in the recursion, which further intensifies memory operations.
@@ 5.1 Baseline CPU performance
Figure 3 presents a comparison of the CPU version of Pepper and the parton-level event generator Comix 666We use a non-default mode for Comix, splitting the process into Monte-Carlo channels grouped by gluons, valence and sea quarks. This is similar to Pepper’s strategy and increases the efficiency of the event generation. [83]. We measure the rate at which unweighted events are generated and stored for and production at the LHC. The center-of-mass energy of the collision is set to TeV, and the partons are required to fulfill , and . For production, an additional cut of is applied to the invariant mass of the dilepton system. For simplicity, the renormalization and factorization scales are fixed to in production, and to in production. For better comparability of the measurements, we show Comix both in combination with its recursive phase space generator [83] and with the modified basic Chili integrator [20] that is also used by Pepper, cf. Sec. 2 for details. The raw data for Fig. 3 are listed in App. A.
@@ 5.2 Performance on different hardware
Using a portability framework like Kokkos in the Pepper event generator is a novel feature for a production-ready parton-level event generator, and for much of the HEP software stack in general. We have therefore tested the performance of the Kokkos Pepper variant against both a native CUDA and a native single-threaded CPU implementation, and looked at results on an A100 GPU and an Intel Core i3-8300 CPU at 3.70 GHz CPU. When comparing the event throughput of the Kokkos and native CUDA implementations on the A100 GPU, we find that the performance of the Kokkos variant agrees within a factor of two, with the performance gap disappearing as computational complexity increases, i.e. with increasing process multiplicity. Generally, the native algorithm outperforms the Kokkos version. This result establishes that using a portability framework is possible without compromising significantly on performance when using a GPU, while at the same time giving access to a much wider range of architectures and computing paradigms. With that, we only show Kokkos result in the following. We note, however, that our current Kokkos implementation on the CPU does not utilize vectorization capabilities, while our native C++ implementation does so with the help of the VCL library [92], which provides significant speedups on the CPU. Because the focus of the analysis here is on portability, these improvements are not yet included in Figs. 4 and 5. We expect to provide a Kokkos implementation with vectorization capabilities in the future. For a more in-depth discussion of CPU vectorization of Pepper, see App. E.
@@ Appendix C Parton-level validation
The comparison results for the two Z+jets observables are shown in Fig. 8. For details on the presentation and the observables, we refer to Sec. 4. The agreement is again quantified using a Kolmogorov–Smirnov test. The -values for each jet multiplicity can be found on the plot. They are all greater than our chosen confidence level of , i.e. we do not reject the null hypothesis of the two underlying distributions being identical.
Finally, the comparison results for the two +jets observables are shown in Fig. 9. Again, the results of a Kolmogorov–Smirnov test are quoted on the plot, with all -values being greater than 0.05. Thus, also in this case, we do not reject the null hypothesis.
@@ Appendix D Data layouts and CPU performance
The two approaches yield different caching opportunities.777 You might also expect them to yield different degrees of CPU auto-vectorization. However, to achieve good performance on the GPU we had to fuse most kernels of the matrix element recursion into a larger recursion kernel, to avoid the overhead of starting a large number of very small kernels on the GPU. This means that the loop over events is a few layers removed from the inner-most loop in this case, which is of course also the most compute-intense part of the code. This seems to prevent auto-vectorization almost entirely. We study two ways to CPU-vectorize the code explicitly in App. E. None of these play a role in the current discussion though, as we have disabled the explicit vectorization for all results obtained in App. D. In Fig. 10, we compare their performance on an Apple M2 Pro chip for different event batch sizes for the jets process. The performance is measured in events per hour, normalized to the SoA event rate for a batch size of one. We find small performance improvements of about when increasing the batch size to 10…100 events, with the SoA performance being within a few percent with the AoS performance, indicating that good cache efficiency is achieved for both data layouts.
@@
Appendix E CPU vectorization
As mentioned in Sec. 5.2, the native C++ implementation of Pepper does use some explicit CPU vectorization for certain calculations. Currently, it does so by implementing real and complex four momentum classes with the help of the Vector Class Library (VCL) [92], such that all the arithmetic operations involving four momenta are compiled using vector intrinsics of the CPU (if supported by VCL). The advantage is that this is a drop-in replacement: One can simply replace the four momentum classes, and any code using these classes immediately profit from the acceleration, without any change. However, there are two disadvantages. One is that the maximum degree of parallelization is limited by the size of the objects. E.g. for AVX-512, where CPU vectors are 512 bits wide, only operations including complex four momenta (which consist of 8 64-bit numbers) would fully utilize the hardware. The second disadvantage is that only simple operations, i.e. addition, subtraction and multiplication by a scalar, are fast. Other operations such as scalar products, that need to take a sum over all components, are only expected to achieve medium efficiency [92]. We can test the performance by measuring the evaluation time of a compute kernel with non-trivial arithmetics and a sizable contribution to the overall runtime. A perfect example is the three gluon vertex kernel, which in particular calculates the complex four-component current
Here, the are complex four-component currents and the are real four momenta, and the subscripts label the three outgoing particles of the vertex. This mixes real and complex vectors and simple component-wise and component-mixing operations. We choose to measure the evaluation time in production on an Intel Xeon Gold 6430 with its support for AVX-512 intrinsics. For 9600 weighted events, the evaluation of the three gluon vertexes requires 4.8 (15.1) seconds with (without) VCL accelerated four momentum classes enabled. Thus, a speed-up of about a factor of three is observed, which falls significantly short of the theoretical factor of eight for double precision arithmetics with 512 bit wide CPU vectorization.
The implemented method discussed so far vectorizes calculations for each individual event. Given the SoA data layout of Pepper, see App. D, it is straightforward to study a different way to vectorize the code, namely to vectorize calculations across several events. This eliminates both disadvantages discussed above: It scales arbitrarily with the size of the CPU vectors, as long as we set the event batch size to some integer multiple of the CPU vector size (measured in double-precision lanes, i.e. multiples of 64 bit). Also, all operations are fully parallel, as the calculations for the events are completely independent from one another. It can therefore be expected that more significant speed-ups can be achieved. Ideally, if the overhead is negligible, the speed-up is equal to the number of CPU double-precision lanes. As a preliminary test, we port the evaluation of the three gluon vertex to use vector intrinsics via the Highway library [103], which supports intrinsics on most common platforms.
We again measure the time for the evaluation of the three gluon vertices in production, on the same Intel Xeon Gold 6430 machine, which supports AVX-512 intrinsics (8 lanes), but also AVX2 intrinsics (4 lanes) and SSE4 intrinsics (2 lanes), such that we can study the scaling with the number of lanes. In addition, we measure it for an Apple M2 Pro chip with ARM NEON intrinsics (2 lanes), to establish the portability advantage of the Highway library as compared to VCL, which does not support ARM and does not plan to do so [92]. To calculate a speed-up, we divide by the time needed on the respective chip when disabling Highway’s use of intrinsics (the operators are then defined by standard C++ code). Note that each time measurement is repeated three times, and then the average is used. We only find small variations across the repetitions which are irrelevant for the present discussion.
Figure 11 shows the achieved speed-ups versus the number of lanes for the three gluon vertex evaluation time for the various vector intrinsics used. While the speed-up is nearly ideal in the 2-lane case, we find speed-ups of 3.8 and 6.9 for the 4- and 8-lane cases, respectively, indicating that overheads become a relevant factor, possibly exacerbated by a reduction of the clock-speed in the AVX-512 case [104]. However, we still get close to the ideal speed-up in all tests and therefore deem this a successful demonstration that Pepper can be vectorized in this manner, with only minor modifications that can be implemented on a kernel-to-kernel basis. The only drawback so far is that in our initial implementation for this demonstration, we lose readability and maintainability of the kernel code because we decompose the calculation of the four currents into its 4 real and 4 imaginary components instead of relying on operators that work directly with the complex currents and four momenta, thus losing the encapsulation of the details of the calculation. We leave it to future work and applications to determine if this is a compromise worth making, or to find a way to wrap the operations in such a way as to sufficiently retain the expressiveness of the kernel code, while still achieving comparable speed-ups.